OpenStudy (anonymous):

How do you find restrictions on the variable?? ALGEBRA 2

OpenStudy (anonymous):

\[n^4-11n^2+30/n^4-7n^2+10\]

OpenStudy (anonymous):

i simplified it and got \[n^2-6/n^2-2\] but i don't know how to find the restrictions?

OpenStudy (anonymous):

Restrictions are where the denominator is equal to 0. What was the denominator after you factored, but before you cancelled?

OpenStudy (anonymous):

\[(n^2-2) (n^2-5)\]

OpenStudy (anonymous):

so the restrictions are 2 and 5?

OpenStudy (anonymous):

ok so you just need to set that equal to 0 and solve for x.

OpenStudy (anonymous):

n² - 2 = 0 and n² - 5 = 0

OpenStudy (anonymous):

right so i would get 2 and 5 positive but i should get the square root?

OpenStudy (anonymous):

You should get positive and negative square roots of 2 and 5

OpenStudy (anonymous):

or is it \[\sqrt{2} and \sqrt{5}\]

OpenStudy (anonymous):

\[x=\pm \sqrt{2},\pm \sqrt{5}\]

OpenStudy (anonymous):

The negatives are also solutions

OpenStudy (anonymous):

okay i get it thanks peach(: